OpenStudy (anonymous):

hi guys im new so im not sure how this works i have a few algebra questions i need help with, if anyone can help

OpenStudy (mathmale):

Please post the questions first, and then ask for help. Just posting them is all you really need to do to get offers of help.

OpenStudy (anonymous):

the first question: Let f(x)=(x^2+3-4) and g(x)=(x+4) find (f)(g) and state the domain then find f/g and state the domain

OpenStudy (mathmale):

Are you sure you've copied this problem correctly? f(x)=(x^2+3-4) I ask this because +3-4=-1 (which seems too elementary to be part of a serious question).

OpenStudy (anonymous):

oh im sorry its meant to be f(x)=(x^2+3x-4) and g(x)=(x-4) sorry for the confusion

OpenStudy (anonymous):

after we go over that first one, can you try helping me out with the others? @mathmale

OpenStudy (mathmale):

"fg" signifies the product of functions f and g. 2 * 5 is the product of 2 and 5 and is 10. Write f(x) and g(x) here. Enclose each one in parentheses. Then multiply the two functions together. Show your work, please.

OpenStudy (mathmale):

f(x) = ? g(x) = ? f(x)*g(x) = ( ? )( ? ) = ??

OpenStudy (anonymous):

wait so do i find f(x) and g(x) individually first? or do i just multiply (x^2+3x-4) to (x+4)

OpenStudy (anonymous):

oh wait i got it!! f(x)=(x+4)(x-1) and g(x)=(x-4) so i get fg=(x+4)^2(x-1) right??

OpenStudy (anonymous):

@mathmale i figure out that one^^^ but i have one more question that i need help with we have the function f(x)=3x+12 a).Determine the inverse of the function and name it g(x) b).Use composite functions to show that these functions are inverses. c). evaluate f(g(-2)) and explain what the domain is. i only need help with b and c, i figured out that the inverse is g(x)=(x-12)/3